OpenStudy (anonymous):

What is the exact value of 4q^6/ sqrt q^21? Simplify if possible.

OpenStudy (anonymous):

\[\large \frac{4q^6}{\sqrt{q^{21}}}=\frac{4q^6}{\sqrt{q^{20}q^1}}=\frac{4q^6}{q^{10}\sqrt{q}}=\frac{4}{q^4 \sqrt{q}} =\frac{4\sqrt q}{q^5}\]